# How Memory Management Impacts LLM Agents: An Empirical Study of Experience-Following Behavior

### v1

@@ 6.2 Memory Management with Resource Constraints
Figure 7: Comparison of unlimited memory size versus a limited memory size of 100 records for strict and coarse selective addition with combined deletion.
We also investigate a scenario where the memory capacity is fixed, for example, to its initial size of 100 records for EHRAgent. Under this constraint, we modify the combined deletion policy so that after each task execution, it first performs periodical-based deletion and then removes only the record with the least average utility (rather than all records below the utility threshold) if the memory still exceeds the limit after additions.
@@ Results
As shown in Figure 6, the memory management policies achieve high performance under strict capacity constraints compared with the fixed-memory variant. By selectively retaining only the most relevant and high-quality records, the agent utilizes limited storage efficiently. These results indicate that effective memory management choices can still improve the long-term agent performance in resource-limited settings. We also study the relationship between the size of the constrained memory and the performance on AgentDriver, as shown in Figure 18 in Appendix C.4, which demonstrates a gradually converging performance when using a strict utility evaluator. This also suggests that naive, unbounded memory growth is unnecessary.
@@ 7 Conclusion
This paper investigates memory management in general LLM agents through addition and deletion. We identify the experience-following phenomenon and its challenges: error propagation and misaligned experience replay. Experiments show that combining deletion with selective addition, guided by a strong evaluator, effectively mitigates these issues. The framework’s robustness and adaptability are further confirmed with two challenge scenarios.
@@ A.1 Agent details and functionality
Agent Name Task Input Output Retrieve Feature #Experiences
EHRAgents EHR tasks Task text query Code Text embedding 4
AgentDriver Autonomous Driving Vehicle state data Predicted trajectory Ego state, goal and history trajectory 1
CIC-IoT Agent IoT intrusion detection (w/ 34 attack type) IoT packet data Reasoning and attack type prediction IoT environment features 3
A.2 Detailed setup on addition, deletion, and challenge scenarios
@@
For the text encoder, we use OpenAI text-embedding-3-large. For the coarse evaluator, we use the built-in LLM-generated termination signal, which reflects if LLM considers the task to be completed. For the strict evaluator, we use string matching to check if the ground-truth is contained in the generated answer. For input similarity, we take the highest cosine similarity between the text embedding of all retrieved experiences and the task. For output similarity, we use pycode_similar package [^3] https://github.com/fyrestone/pycode_similar.git to detect the code plagiarism score between the execution from the retrieved memory with the highest input similarity and current model execution. For periodic deletion, we use a period of 200 and $\alpha=0$. For history-based deletion, we set the minimal deleting frequency to 5 and threshold $\beta$ to 0.3 for the coarse evaluator and 0.7 for the strict evaluator. The combined deletion reuses all hyper-parameters above.
@@
For the retrieval strategy, the original retrieval pipeline proposed by [13] consists of a two-step process: first, selecting the top-3 experiences based on vector similarity, and then prompting an LLM to choose the most relevant one. To improve reproducibility, we simplify this process to a single-step retrieval, selecting only the top-1 experience based on vector similarity. For the coarse evaluator, we employ an LLM to assess the quality of retrieved experiences. The LLM outputs either “yes” or “no” to indicate whether a given experience is a “good” experience or a “bad” one. The prompt used for this LLM-based evaluation is provided in Appendix A.4. For the strict evaluator, we measure the UniAD 3-second average L2 distance between the predicted trajectory and the ground truth. Experiences with an L2 error lower than 2.5 are added to the memory bank. For input similarity, we adhere to the computation method from the original paper. Specifically, we compute the exponential negative L2 distance between the query vector and stored experiences, using the same coefficients as in [13]. For output similarity, we compute the radial basis function (RBF) kernel between the predicted trajectory and the ground-truth trajectory. The similarity score is calculated using the following formula:
[…]
In our experiments, we set $\gamma=1.0$. For periodic deletion, we apply a deletion period of 300 steps with a threshold of $\alpha=1$. For history-based deletion, we set a threshold of 5.0 for the UniAD 3-second average L2 distance. An experience is deleted if it has been retrieved at least 5 times and the mean UniAD 3-second average L2 distance across all retrievals exceeds this threshold. The combined deletion mechanism incorporates all the aforementioned hyperparameters.
@@
For the text encoder, we use OpenAI text-embedding-3-large. For the coarse evaluator, similar to AgentDriver, we employ an LLM to assess the quality of the experiences. The prompt for the CIC-IoT Agent’s LLM evaluator is provided in Appendix A.4. For the strict evaluator, we use string matching to check if the ground-truth is contained in the generated answer. For input similarity, we use a feature-based approach. To measure input similarity, we adopt a feature-based approach. Specifically, we compute the relative change across all features and take the average. For continuous features, the relative change between two inputs, $\mathbf{input}_{1}$ and $\mathbf{input}_{2}$, for feature $f_{i}$ is computed as:
@@
For output similarity, we calculate the embedding similarity. For periodic deletion, we set a cycle of 300 steps, and $\alpha=1$. For history-based deletion, we set the experience quality threshold $\beta=0.5$. If an experience is retrieved 5 or more times and the average score of all test samples when using this experience falls below 0.5, the experience is removed. The combined deletion strategy reuses all the hyperparameters mentioned above.
@@ A.3 CIC-IoT Agent design
Below is the user prompt used in CIC-IoT Agent to incorporate IoT packet data and all possible attack types.
User Prompt Based on the following features, determine the most likely attack type from the list below: Flow duration [description: Duration of the packet’s flow]: {flow_duration} Header Length [description: Header Length]: {Header_Length} bytes Protocol Type [description: IP, UDP, TCP, IGMP, ICMP, Unknown (Integers)]: {Protocol_Type} Duration [description: Time-to-Live (ttl)]: {Duration} Rate [description: Rate of packet transmission in a flow]: {Rate} Srate [description: Rate of outbound packets transmission in a flow]: {Srate} Drate [description: Rate of inbound packets transmission in a flow]: {Drate} Number of FIN flags [description: FIN flag value]: {fin_flag_number} Number of SYN flags [description: SYN flag value]: {syn_flag_number} Number of RST flags [description: RST flag value]: {rst_flag_number} Number of PSH flags [description: PSH flag value]: {psh_flag_number} Number of ACK flags [description: ACK flag value]: {ack_flag_number} Number of ECE flags [description: ECE flag value]: {ece_flag_number} Number of CWR flags [description: CWR flag value]: {cwr_flag_number} Number of ACK packets [description: Number of packets with ACK flag set in the same flow]: {ack_count} Number of SYN packets [description: Number of packets with SYN flag set in the same flow]: {syn_count} Number of FIN packets [description: Number of packets with FIN flag set in the same flow]: {fin_count} Number of URG packets [description: Number of packets with URG flag set in the same flow]: {urg_count} Number of RST packets [description: Number of packets with RST flag set in the same flow]: {rst_count} HTTP traffic flag [description: Indicates if the application layer protocol is HTTP]: {HTTP} HTTPS traffic flag [description: Indicates if the application layer protocol is HTTPS]: {HTTPS} DNS traffic flag [description: Indicates if the application layer protocol is DNS]: {DNS} (Continued on next page…)
@@
Telnet traffic flag [description: Indicates if the application layer protocol is Telnet]: {Telnet}
SMTP traffic flag [description: Indicates if the application layer protocol is SMTP]: {SMTP} SSH traffic flag [description: Indicates if the application layer protocol is SSH]: {SSH} IRC traffic flag [description: Indicates if the application layer protocol is IRC]: {IRC} TCP traffic flag [description: Indicates if the transport layer protocol is TCP]: {TCP} UDP traffic flag [description: Indicates if the transport layer protocol is UDP]: {UDP} DHCP traffic flag [description: Indicates if the application layer protocol is DHCP]: {DHCP} ARP traffic flag [description: Indicates if the link layer protocol is ARP]: {ARP} ICMP traffic flag [description: Indicates if the network layer protocol is ICMP]: {ICMP} IPv4 traffic flag [description: Indicates if the network layer protocol is IP]: {IPv} LLC traffic flag [description: Indicates if the link layer protocol is LLC]: {LLC} Total sum of feature values [description: Summation of packets’ lengths in the flow]: {Tot_sum} Minimum value [description: Minimum packet length in the flow]: {Min} Maximum value [description: Maximum packet length in the flow]: {Max} Average value [description: Average packet length in the flow]: {AVG} Standard deviation [description: Standard deviation of packet length in the flow]: {Std} Total size of the flow [description: Packet’s length]: {Tot_size} bytes Inter-arrival time (milliseconds) [description: The time difference with the previous packet]: {IAT} Number of packets or flows [description: The number of packets in the flow]: {Number} Magnitude of the flow [description: Average of the lengths of incoming packets in the flow + average of the lengths of outgoing packets in the flow]: {Magnitude} Radius of the flow [description: Variance of the lengths of incoming packets in the flow + variance of the lengths of outgoing packets in the flow]: {Radius} Covariance of the flow [description: Covariance of the lengths of incoming and outgoing packets]: {Covariance} Variance of the flow [description: Variance of the lengths of incoming packets in the flow / variance of the lengths of outgoing packets in the flow]: {Variance} Weight of the flow [description: Number of incoming packets / number of outgoing packets]: {Weight} ### Attack Types: [’Backdoor_Malware’, ’BrowserHijacking’, ’DDoS-UDP_Flood’, ’DDoS-SynonymousIP_Flood’, ’Uploading_Attack’, ’Mirai-udpplain’, ’DDoS-SlowLoris’, ’DDoS-ICMP_Flood’, ’Recon-OSScan’, ’DoS-HTTP_Flood’, ’DNS_Spoofing’, ’BenignTraffic’, ’DoS-UDP_Flood’, ’DictionaryBruteForce’, ’DoS-SYN_Flood’, ’Recon-PortScan’, ’Recon-HostDiscovery’, ’DDoS-PSHACK_Flood’, ’Mirai-greeth_flood’, ’VulnerabilityScan’, ’DDoS-RSTFINFlood’, ’Recon-PingSweep’, ’DDoS-UDP_Fragmentation’, ’DDoS-HTTP_Flood’, ’XSS’, ’Mirai-greip_flood’, ’DoS-TCP_Flood’, ’MITM-ArpSpoofing’, ’DDoS-ACK_Fragmentation’, ’CommandInjection’, ’SqlInjection’, ’DDoS-ICMP_Fragmentation’, ’DDoS-TCP_Flood’, ’DDoS-SYN_Flood’]

### v2

@@ 5.2 Memory Management with Resource Constraints
We also investigate a scenario where the memory capacity is fixed, for example, to its initial size of 100 records for EHRAgent and 180 records for AgentDriver. Under this constraint, we modify the combined deletion policy so that after each task execution, it first performs periodical-based deletion and then removes only the record with the least average utility (rather than all records below the utility threshold) if the memory still exceeds the limit after additions.
@@ Results
As shown in Figure 17 in Appendix B.6, the memory management policies achieve high performance under strict capacity constraints compared with the fixed-memory variant. By selectively retaining only the most relevant and high-quality records, the agent utilizes limited storage efficiently. These results indicate that effective memory management choices can still improve the long-term agent performance in resource-limited settings. We also study the relationship between the size of the constrained memory and the performance on AgentDriver, as shown in Figure 18 in Appendix B.6, which demonstrates a gradually converging performance when using a strict utility evaluator. This also suggests that naive, unbounded memory growth is unnecessary.
@@ 6 Conclusion
This paper studies memory management in LLM agents through addition and deletion. We identify the experience-following phenomenon and reveal two key challenges—error propagation and misaligned experience replay. Experiments demonstrate that evaluator reliability is critical and that incorporating evaluator signals is essential for effective memory management.
Limitations
One limitation of this work lies in the scope of memory management considered. To build a generalizable understanding of memory dynamics over time, this study focuses on two fundamental operations—memory addition and memory deletion—while omitting more complex and less generalizable mechanisms such as structural transformation Zeng et al. (2024), merging Liu et al. (2023), summarization Pan et al. (2025), and reflection Shinn et al. (2024). Consequently, when extending our findings to systems that integrate these advanced updating mechanisms or operate with different agent architectures, additional fine-grained analyses will be helpful.
Another limitation concerns the lack of theoretical guarantees. Our findings are primarily based on extensive empirical analyses, which, although comprehensive, do not provide formal theoretical proofs. Given the inherent complexity of agentic systems, deriving such guarantees is often infeasible. Nonetheless, we believe that the proposed framework offers a solid foundation for future theoretical investigation, particularly through controlled experimentation in our synthetic RegAgent environment.
@@ A.1 Agent details and functionality
Agent Name Task Input Output Retrieve #Experiences
Feature
RegAgent Regression prediction Input vector Output number Input vector 6
EHRAgents EHR tasks Task Code Text 4
text query embedding
AgentDriver Autonomous Vehicle Predicted trajectory Ego state, goal, 1
Driving state data history trajectory
CIC-IoT Agent IoT Traffic Detection IoT packet Reasoning and prediction IoT features 3
(8 categories) data attack type
RegAgents
RegAgent serves as a synthetic and controlled environment to study the effects of memory management on long-term agent performance. We first generate an implicit weight vector $w\in\mathbb{R}^{6}$, and the task is to predict the correct output $y=w^{\top}x$ by retrieving the six most relevant demonstrations from memory. Specifically, we generate 100 input vectors $x$ sampled from three Gaussian distributions $N(\mu,1)$ with $\mu\in{-0.5,0,0.5}$, each of six dimensions, and compute $y=w^{\top}x+\epsilon$, where $\epsilon$ is bounded noise within [1,1][-1,1]. These pairs form the initial memory bank. During actual execution, we collect 4000 input–output pairs from the same three distributions. A prediction $\hat{y}$ is considered successful if $|\hat{y}-y|\leq 1$, and we use the success rate as the primary performance metric.
For retrieval, input similarity is measured by cosine similarity between input vectors, while output similarity is defined as $\text{output\_similarity}=\exp(-\gamma|\mathbf{x}_{1}-\mathbf{x}_{2}|^{2}),$ where $\gamma=1.0$.
This task is highly controllable and well-suited for our study, as it (1) relies solely on the retrieved demonstrations, (2) allows the agent to either reason or directly mimic the demonstrations to produce outputs, and (3) enables clear definitions of error and addition criteria.
For evaluation, the strict evaluator applies a threshold of 1.0 on the absolute error between prediction and ground truth. Coarse evaluators use thresholds of 1.6 (C1), 1.4 (C2), and 1.2 (C3), respectively. Periodic deletion is performed with 500 steps with $\alpha=0$. History-based deletion is applied with a minimum deletion frequency of 5 and a threshold $\beta=0.5$ across all setups. The combined deletion strategy reuses all hyperparameters above. We also list the task prompt in Appendix A.2
@@
Following EHRAgent Shi et al. (2024), our agent is a code-generation system that enables clinicians to interact with electronic health records (EHRs) through natural language queries. We adopt the MIMIC-III dataset Johnson et al. (2016) for evaluation. After filtering out duplicated and unanswerable entries, we obtain 2,392 tasks in total. During each test query, the agent retrieves four past experiences from an initial memory bank containing 100 records, based on maximum cosine similarity computed using a general text embedding model. We report task performance in terms of accuracy (ACC).
For our implementation, we use OpenAI text-embedding-3-large as the text encoder. For the coarse evaluator, we use the built-in LLM-generated termination signal, which reflects if LLM considers the task to be completed. For the strict evaluator, we use exact match to determine whether the generated answer is correct. For input similarity, we take the highest cosine similarity between the text embedding of all retrieved experiences and the task. For output similarity, we use the pycode_similar package [^3] https://github.com/fyrestone/pycode_similar.git to detect the code plagiarism score between the execution from the retrieved memory with the highest input similarity and the current model execution. For periodic deletion, we use a period of 200 and $\alpha=0$. For history-based deletion, we set the minimal deleting frequency to 5 and the threshold $\beta$ to 0.3 for the coarse evaluator using GPT-4o-mini and GPT-4.1-mini, and 0.7 for the strict and GPT-4.1-mini fine-tuned evaluator. The combined deletion reuses all hyperparameters above.
@@
Following AgentDriver Mao et al. (2024), we adopt an LLM-based autonomous driving agent that integrates common sense and experiential knowledge into memory records. We use the nuScenes dataset Caesar et al. (2020) and randomly sample 2,000 test cases from its test split. The initial memory contains 180 experiences randomly drawn from the nuScenes training set.
For the retrieval strategy, the original retrieval pipeline proposed by Mao et al. (2024) consists of a two-step process: first, selecting the top-3 experiences based on vector similarity, and then prompting an LLM to choose the most relevant one. To improve reproducibility, we simplify this process to a single-step retrieval, selecting only the top-1 experience based on vector similarity. For the coarse evaluator, we employ an LLM to assess the quality of retrieved experiences. The LLM outputs either “yes” or “no” to indicate whether a given experience is a “good” experience or a “bad” one. The prompt used for this LLM-based evaluation is provided in Appendix A.4. For the strict evaluator, we measure the UniAD 3-second average L2 distance between the predicted trajectory and the ground truth. Experiences with an L2 error lower than 2.5 are added to the memory bank. For input similarity, we adhere to the computation method from the original paper. Specifically, we compute the exponential negative L2 distance between the query vector and stored experiences, using the same coefficients as in Mao et al. (2024). For output similarity, we compute the radial basis function (RBF) kernel between the predicted trajectory and the ground-truth trajectory. The similarity score is calculated with the same formula in RegAgent:
$\text{output\_similarity}=\exp\left(-\gamma\|\mathbf{v_{1}}-\mathbf{v_{2}}\|^{2}\right).$
In our experiments, we set $\gamma=1.0$. For periodic deletion, we apply a deletion period of 500 steps with a threshold of $\alpha=0$. For history-based deletion, we set a threshold of 5.0 for the UniAD 3-second average L2 distance. An experience is deleted if it has been retrieved at least 3 times and the mean UniAD 3-second average L2 distance across all retrievals exceeds this threshold. Coarse deletion leverages the predicted success rate from the coarse evaluator and deletes a record if the accumulated success rate is less than 0.5. The combined deletion mechanism incorporates all the aforementioned hyperparameters.
@@
Following the setup in the CIC-IoT benchmark Neto et al. (2023), we implement a network traffic detection agent designed to predict traffic types from IoT packet features. The original CIC-IoT dataset contains 34 attack classes. Since our implementation relies solely on single-flow features, we exclude categories that cannot be distinguished based on such features. For example, we retain DDoS- HTTP_Flood while removing DoS- HTTP_Flood due to their high similarity under single-flow representations. After filtering, we retain 8 representative classes that remain distinguishable using single-flow features, and construct each test query by formatting these features into a predefined prompt template. We evaluate CIC-IoT Agent on the CIC-IoT dataset, randomly sampling 1,000 test cases from its training split. During each execution, the agent retrieves three past experiences using a feature-based similarity approach. The initial memory bank is initialized with 100 synthetic records generated by GPT-4o-mini from a disjoint training subset. We report task performance in terms of accuracy (ACC).
For the text encoder, we use OpenAI text-embedding-3-large. For the coarse evaluator, similar to AgentDriver, we employ an LLM to assess the quality of the experiences. The prompt for the CIC-IoT Agent’s LLM evaluator is provided in Appendix A.4. For the strict evaluator, we use string matching to check if the ground-truth is contained in the generated answer. For input similarity, we use a feature-based approach. Specifically, we compute the relative change across all features and take the average. For continuous features, the relative change between two inputs, $\mathbf{input}_{1}$ and $\mathbf{input}_{2}$, for feature $f_{i}$ is computed as:
@@
For output similarity, we calculate the embedding similarity. For periodic deletion, we set a cycle of 500 steps, and $\alpha=1$. For history-based deletion, we set the experience quality threshold $\beta=0.7$. If an experience is retrieved 3 or more times and the average score of all test samples when using this experience falls below 0.7, the experience is removed. The combined deletion strategy reuses all the hyperparameters mentioned above.
A.2 RegAgent design
Below is the user prompt used for the RegAgent task.
User Prompt You are given a 6-dimensional input vector x. Predict $y=w^{T}x$ with an unknown w. You will see K demonstrations of (input, guess) pairs that use the same w but may contain noise in all demonstrations. You need to strictly follow the output content and format of the demonstrations, which is Guess: boxed{{<number>}} without any other text. Demonstrations (K=k): {demonstrations} Now solve for the new input. Input: {x} Guess:
@@ A.3 CIC-IoT Agent design
Below is the user prompt used in CIC-IoT Agent to incorporate IoT packet data and all possible traffic types.
User Prompt Based on the following features, determine the most likely traffic type from the list below: * Pay special attention to cross-field consistency checks. * Do not be misled by a single feature if it conflicts with others. * Note: Your reasoning should be based on all features, not on any single field. You are allowed to select only one traffic type as your answer. If you choose more than one, your answer will be marked as incorrect. Required output format: ANALYSIS: {your reasoning here, including key features and justification} ANSWER: {traffic_type} (Continued on next page…)
User Prompt Flow duration [description: Duration of the packet’s flow]: {flow_duration} Header Length [description: Header Length]: {Header_Length} bytes Protocol Type [description: IP, UDP, TCP, IGMP, ICMP, Unknown (Integers)]: {Protocol_Type} Duration [description: Time-to-Live (ttl)]: {Duration} Rate [description: Rate of packet transmission in a flow]: {Rate} Srate [description: Rate of outbound packets transmission in a flow]: {Srate} Drate [description: Rate of inbound packets transmission in a flow]: {Drate} Number of FIN flags [description: FIN flag value]: {fin_flag_number} Number of SYN flags [description: SYN flag value]: {syn_flag_number} Number of RST flags [description: RST flag value]: {rst_flag_number} Number of PSH flags [description: PSH flag value]: {psh_flag_number} Number of ACK flags [description: ACK flag value]: {ack_flag_number} Number of ECE flags [description: ECE flag value]: {ece_flag_number} Number of CWR flags [description: CWR flag value]: {cwr_flag_number} Number of ACK packets [description: Number of packets with ACK flag set in the same flow]: {ack_count} Number of SYN packets [description: Number of packets with SYN flag set in the same flow]: {syn_count} Number of FIN packets [description: Number of packets with FIN flag set in the same flow]: {fin_count} Number of URG packets [description: Number of packets with URG flag set in the same flow]: {urg_count} Number of RST packets [description: Number of packets with RST flag set in the same flow]: {rst_count} HTTP traffic flag [description: Indicates if the application layer protocol is HTTP]: {HTTP} (Continued on next page…)
@@
HTTPS traffic flag [description: Indicates if the application layer protocol is HTTPS]: {HTTPS}
DNS traffic flag [description: Indicates if the application layer protocol is DNS]: {DNS} Telnet traffic flag [description: Indicates if the application layer protocol is Telnet]: {Telnet} SMTP traffic flag [description: Indicates if the application layer protocol is SMTP]: {SMTP} SSH traffic flag [description: Indicates if the application layer protocol is SSH]: {SSH} IRC traffic flag [description: Indicates if the application layer protocol is IRC]: {IRC} TCP traffic flag [description: Indicates if the transport layer protocol is TCP]: {TCP} UDP traffic flag [description: Indicates if the transport layer protocol is UDP]: {UDP} DHCP traffic flag [description: Indicates if the application layer protocol is DHCP]: {DHCP} ARP traffic flag [description: Indicates if the link layer protocol is ARP]: {ARP} ICMP traffic flag [description: Indicates if the network layer protocol is ICMP]: {ICMP} IPv4 traffic flag [description: Indicates if the network layer protocol is IP]: {IPv} LLC traffic flag [description: Indicates if the link layer protocol is LLC]: {LLC} Total sum of feature values [description: Summation of packets’ lengths in the flow]: {Tot_sum} Minimum value [description: Minimum packet length in the flow]: {Min} (Continued on next page…)
User Prompt (Continuation from previous page…)
Maximum value [description: Maximum packet length in the flow]: {Max}
Average value [description: Average packet length in the flow]: {AVG} Standard deviation [description: Standard deviation of packet length in the flow]: {Std} Total size of the flow [description: Packet’s length]: {Tot_size} bytes Inter-arrival time (milliseconds) [description: The time difference with the previous packet]: {IAT} Number of packets or flows [description: The number of packets in the flow]: {Number} Magnitude of the flow [description: Average of the lengths of incoming packets in the flow + average of the lengths of outgoing packets in the flow]: {Magnitude} Radius of the flow [description: Variance of the lengths of incoming packets in the flow + variance of the lengths of outgoing packets in the flow]: {Radius} Covariance of the flow [description: Covariance of the lengths of incoming and outgoing packets]: {Covariance} Variance of the flow [description: Variance of the lengths of incoming packets in the flow / variance of the lengths of outgoing packets in the flow]: {Variance} Weight of the flow [description: Number of incoming packets / number of outgoing packets]: {Weight} ### Traffic Types: [’DDoS-ICMP_Flood’, ’DDoS-UDP_Flood’, ’DDoS-TCP_Flood’, ’DDoS-SYN_Flood’, ’DDoS-PSHACK_Flood’, ’DDoS-RSTFINFlood’, ’DDoS-HTTP_Flood’, ’BenignTraffic’]